OpenStudy (anonymous):

Evaluate 2xy - m when x = -4, y = 3, and m = 6

OpenStudy (the_fizicx99):

2(-4)(3) - 6

OpenStudy (anonymous):

2(-4)(3) - (6) 2 * -4 = -8 * 3 = -24 -24 - 6 = -30